### v1

# Contextual modulation of language comprehension in a dynamic neural model of lexical meaning

Abstract
We propose and computationally implement a dynamic neural model of lexical meaning, and experimentally test its behavioral predictions. We demonstrate the architecture and behavior of the model using as a test case the English lexical item have, focusing on its polysemous use. In the model, have maps to a semantic space defined by two continuous conceptual dimensions, connectedness and control asymmetry, previously proposed to parameterize the conceptual system for language. The mapping is modeled as coupling between a neural node representing the lexical item and neural fields representing the conceptual dimensions. While lexical knowledge is modeled as a stable coupling pattern, real-time lexical meaning retrieval is modeled as the motion of neural activation patterns between metastable states corresponding to semantic interpretations or readings. Model simulations capture two previously reported empirical observations: (1) contextual modulation of lexical semantic interpretation, and (2) individual variation in the magnitude of this modulation. Simulations also generate a novel prediction that the by-trial relationship between sentence reading time and acceptability should be contextually modulated. An experiment combining self-paced reading and acceptability judgments replicates previous results and confirms the new model prediction. Altogether, results support a novel perspective on lexical polysemy: that the many related meanings of a word are metastable neural activation states that arise from the nonlinear dynamics of neural populations governing interpretation on continuous semantic dimensions.
keywords:
lexical semantics , dynamic field theory , language comprehension[inst1]organization=Linguistics Department, Yale University,addressline=370 Temple St., city=New Haven, state=Connecticut, postcode=06511, country=USA
We propose a model of lexical meaning in the framework of Dynamic Field Theory (DFT)
Model simulations capture known sentence comprehension effects and generate a novel prediction
Sentence reading experiment replicates previous results and confirms new model prediction
We argue that polysemy arises from metastability in neural dynamics on continuous dimensions
Contents
1 Introduction
In language comprehension, linguistic forms evoke interpretations of meaning. A basic unit of linguistic form is the lexical item, the systematic relationship between phonetic/phonological, morphological, syntactic, and semantic information (Anderson, 1992; Jackendoff, 1975). It is a well-known observation that the specific interpretation or reading evoked by a lexical item can vary depending on context. Polysemy refers to the situation whereby a lexical item offers more than one interpretation yet those possibilities are conceptually related. For example, the lexical item book can be described as polysemous since its possible interpretations include (at least) both a physical object (e.g., heavy book) and a collection of information represented by an object (e.g., enjoyable book) (e.g., Brugman, 1988; Deane, 1988; Lakoff, 1990; Pustejovsky, 1995; Vicente, 2018).111Polysemy contrasts with homophony, the situation where the possible interpretations appear semantically unrelated, e.g., river bank vs. savings bank. Here we investigate the conceptual basis of the constrained semantic variability observed in lexical polysemy. To this end, we examine a possible source of this variability, propose a neuro-computational implementation, and investigate the behavioral consequences during sentence comprehension. Under our proposal, the many-to-one relation between meaning and form which is salient in cases of lexical polysemy can be seen as the norm rather than the exception, with differences primarily in the magnitude of variability and the relatedness of the possible interpretations.
1.1 English ‘have’
Our test case is the English lexical item have. have is typically interpreted as a relation of possession between the referents of its arguments, as in 1.1 where the possession is alienable, or as in 1.1 where the possession is inalienable.
. The linguistics professor has a motorcycle. (alienable possession)
. The oak tree has a healthy trunk. (inalienable possession)
When subjects are presented with less frequent argument combinations, as in 1.1 below, two types of responses are reported: (a) the inanimate subject “the oak tree” leads to an inalienable possession interpretation, but this conflicts with the fact that the object “the motorcycle” does not plausibly enter into a part-whole relation with “the oak tree”, leading to a decrease in acceptability; (b) the referent of the subject is anthropomorphized, e.g., “the person dressed as a tree…”, in order to maintain an alienable possession interpretation (Zhang et al., 2018; Zhang, 2021). These facts are consistent with a bias for have to evoke a possession interpretation.
. #The oak tree has a motorcycle.222“#” signifies that the sentence, while grammatical, is judged less felicitous without additional context.
But the polysemy reportoire of have is larger than these two readings. have can also evoke an interpretation of coincidental location. This would render 1.1 above with an interpretation that the motorcycle is in coincidental spatial proximity to (e.g., next to, under, below, above) the oak tree, a situation that we will refer to as adjacency (e.g., Myler, 2016; Zhang et al., 2022). The tendency for have to evoke a possession reading is very strong, but not indefeasible. One standard way to bring out an adjacency reading from have is to add a prepositional phrase which makes explicit the spatial relation, as in 1.1. Not surprisingly, presence of the modifier “next to it” improves acceptability ratings relative to sentences like 1.1 (Zhang et al., 2018, 2022).
. The oak tree has a motorcycle next to it.
This said, presence of an explicit modifier is not the only way to make an adjacency reading salient. A bias for an adjacency interpretation can also be induced by the preceding context. Specifically, when a preceding sentence evokes an adjacency interpretation, as in 1.1, then sentences like 1.1 are more likely to trigger an adjacency interpretation as well as receive higher acceptability ratings (Zhang et al., 2018, 2022), as seen in 1.1.
. The pine tree has a car next to it and the oak tree has a motorcycle.
This linguistic and behavioral pattern suggests a unified polysemy-based analysis of English have involving readings raging from a purely coincidental spatial relation to an alienable possession relation to an inalienable possession relation. Crucially, these readings are not disconnected. The distinction between adjacency, alienable possession, and inalienable possession is a matter of degree, not category, not only between these three reading types but also within them. In the case of coincidentality in spatial location, for example, the situation conveyed by “the tree has a motorcycle” is more coincidental than “the tree has a bench”, and that, in turn, is more coincidental than “the tree has a swing”, or “the tree has a nest”. So, with each situation we move towards less coincidentality and, interestingly, towards greater asymmetry between arguments such that the first argument is more and more perceived as possessing the second.
Similar gradability is observable within the alienable and inalienable possession spaces. For alienable possession, “the woman has a car” can refer to a variety of related situations distinguishable by degree of ownership: the woman has a car because she stole it, because she borrowed it, rented it, bought it. Only in the last instance would she “own” the car, yet in all instances the relation between the woman and the car—one of asymmetry in control—renders it one of possession. Finally, for the inalienable possession space, the possibilities range from the peripheral “the woman has hair” to the more inalienable “the woman has a liver”, to the completely inalienable “the woman has a body”. Notably, while there is no coincidentality in these examples, making it more like the alienable possession space, the control relation between arguments—e.g., woman to her hair, woman to her liver, woman to her body—has become less and less asymmetrical, similar to the relation in the coincidental location space (Koch, 2012; Piñango, 2023; Zhang, 2021).
We conclude then that the lexical meaning of have, while constrained, is variable along continuous dimensions. As a result, such meaning is best characterized as a space within which specific readings can obtain. This conclusion in turn raises the question of what the properties of such a continuous space are that give rise to the constrained variability in meaning that have exhibits. We address this question below.
1.2 Continuous conceptual space for lexical meaning
It has been proposed that the readings of have are supported by a conceptual space organized in terms of two continuous cognitive dimensions or metrics: control asymmetry and connectedness (Gärdenfors and Warglien, 2012; Piñango, 2019, 2023; Zhang, 2021). Control asymmetry measures the degree to which two individuals in a situation differ in how much one controls the other. Control asymmetry underpins perception of causality and causal chains. In doing so, it gives rise to intuitions about intentionality and agency (e.g., Carey, 2009; Croft, 2012; Talmy, 1988). It has also been shown to constrain language development (e.g., Klein and Perdue, 1992). An evaluation of high control asymmetry between two participants means that one of the participants can be construed as a controller in the situation, that is, an intentional agent, and not the other way around. A low control asymmetry evaluation means that neither of the participants is construable as controller. In the case of have, ultimate evaluation of control asymmetry depends on the context, including the properties of the participants.
Connectedness measures the degree to which two individuals in a situation are functionally or structurally related to one another. It is from connectedness that intuitions emerge about coherence in the world, e.g., object individuation (e.g., Carey, 2009; Gopnik et al., 2004; Krøjgaard, 2004). Relations of connectedness are built along independently motivated conceptual dimensions—spatial, temporal, informational, or functional (Piñango and Deo, 2016). In the case of have, ultimate evaluation of connectedness again depends on the context and the perceived properties of the argument referents.
In Figure 1, three possible interpretations of have are plotted in the space spanned by the two dimensions of connectedness (horizontal axis) and control asymmetry (vertical axis). A coincidental adjacency interpretation, as in 1.1, corresponds to low connectedness and low control asymmetry, and an alienable possession interpretation, as in 1.1, corresponds to higher connectedness and higher control asymmetry. An inalienable possession interpretation, as in 1.2, corresponds to high connectedness and low control asymmetry.
. The oak tree has a thick trunk.
As seen in Figure 2, when more possible interpretations of have are plotted (including, in addition to the examples given so far, at least containment, control, and kinship), they tend to fall on an inverted “U” shape. This pattern indicates dependencies between the two dimensions which constrain object perception and its linguistic encoding. We return to this point in the context of our model in Section 2.
1.3 Further empirical evidence for continuous semantic dimensions
In addition to the robust observation of gradability between readings of have, empirical support for a model of lexical meaning based on interpretable, continuous dimensions comes from two other sources: (a) trajectories of meaning change over time, and (b) real-time sentence comprehension effects. Regarding (a), over an approximately 200-year period, the postposition kade in the Indo-Aryan language Marathi has shifted gradually from primarily an adjacency reading, to primarily an alienable possession reading, to primarily an inalienable possession reading (Deo, 2015; Zhang, 2021). This trajectory essentially traces the inverted “U” shape in Figure 2 from left to right. This supports the notion that the potential readings of have are organized along these continuous dimensions, rather than forming a countable set of discrete meanings.
Regarding (b), evidence comes from acceptability judgments, self-paced reading, and electroencephalography (EEG) experiments investigating the availability of the coincidental adjacency reading associated with have as a result of preceding context. Zhang et al. (2018) administered an acceptability judgment task with target sentences consisting of have and two inanimate arguments, as in 1.1. Each target sentence was preceded by a context sentence, also consisting of have and two inanimate arguments. The crucial manipulation was the meaning of the context sentence: adjacency, as in 1.1, or inalienable possession, as in 1.3.
. The pine tree has big branches.
When preceded by a context sentence designed to evoke an adjacency interpretation, the target sentence received higher acceptability ratings, relative to when the same sentence was preceded by a sentence with a possession meaning. This improvement in acceptability was attributed to a gradient shift in interpretation towards adjacency due to the influence of the context sentence.
Using a similar stimulus set as in the acceptability judgment task, Zhang et al. (2018) observed decreased reading times for target sentences preceded by an adjacency context sentence, relative to when they were preceded by a possession context sentence. Moreover, EEG recordings indicated an N400 event-related potential (ERP) in the possession condition relative to the adjacency condition (Zhang et al., 2018). These results are consistent with a decreased neurocognitive processing load in the adjacency condition, due to contextual facilitation of an adjacency interpretation of the target sentence.
Finally, the magnitude of contextual modulation of have interpretation correlates at an individual level with an independent measure of context sensitivity, i.e., the Autism-Spectrum Quotient (AQ: Baron-Cohen et al., 2001). Individuals with higher AQ scores (decreased context sensitivity) exhibit a decreased difference in acceptability between conditions (Zhang et al., 2022). The influence of context on interpretation is a gradient rather than categorical phenomenon. Some of the variation in the magnitude of contextual influence can be explained by the AQ.
Taken together, observations from historical change, acceptability judgments, self-paced reading, EEG, and individual variation all point to gradience in the cognitive representation of lexical meaning. Such gradience is captured by a parameterized conceptual space such as the one schematized in Figures 1 and 2. The question that arises now is what are the neurocognitive properties of such a system. To investigate this question we turn to Dynamic Field Theory (DFT: Schöner et al., 2016), a formal framework for understanding the neural activation dynamics underlying continuous cognitive dimensions.
1.4 Dynamic Field Theory (DFT)
In this subsection, we give a brief overview of DFT; a more detailed description of the relevant mathematics is given in Section 2. In DFT, features relevant for cognition are modeled as continuous parameters governed by the activity of populations of neurons. The activity of a neural population over time is described using a differential equation defining a dynamic neural field (DNF: Amari, 1977). DNFs are characterized by point attractor dynamics. This means that, at any given time, the activation pattern in a DNF is attracted to a particular state, i.e., the point attractor state. The location of the point attractor can change over time according to a variety of factors, e.g., inputs to the DNF from sensory surfaces or other DNFs. Usually, when a DNF is not receiving any input, the point attractor corresponds to an inactive or resting state. When a DNF begins to be influenced by input, the point attractor might shift to an active state. In particular, the dynamics of lateral interaction within DNFs allow the formation of “peaks” of activation. Depending on the cognitive dimension being represented by the DNF, an activation peak might correspond to a movement goal, a percept, or another kind of cognitive event. Since change within DNFs is characterized by point attractor dynamics, but the location of the point attractor is itself ever-changing, DNFs can be described as metastable (e.g., Kelso, 2012). Metastability, invoking the interplay between stability and flexibility, is an important property of cognition (Kelso, 1995). Stability offers resistance to the ubiquitous influence of noise, and flexibility allows rapid change under changing cognitive and environmental conditions. In the context of lexical meaning, metastability offers a way to reconcile intuitions of discreteness with evidence for underlying continuity.
DFT originally developed in the context of motor control research, especially in the domains of eye movements (Kopecz and Schöner, 1995) and arm movements (Erlhagen and Schöner, 2002). It has been increasingly applied in other cognitive domains, including, recently, speech and language. DFT models of speech and language have, so far, focused on the neurocognitive representation of phonetic dimensions, e.g., voice onset time (VOT), and the location and degree of constrictions formed by the tongue. Thus, activation peaks correspond to articulatory movement goals. These models have offered novel explanations for a variety of empirical phenomena: effects of auditory perception on verbal response times (Roon and Gafos, 2016), effects of lexical competitors on speech articulation in errors (Stern et al., 2022) and non-errors (Stern and Shaw, 2023a, b), long-term phonological change (Gafos and Kirov, 2009; Shaw and Tang, 2023), and individual differences in phonological representations (Harper, 2021).
1.5 This paper
In this study, we extend neural field dynamics to dimensions of semantic interpretation. While the relevant cognitive dimensions are specific to the meaning system and therefore differ from those of previous DFT models of speech, the basic dynamics of neural activity are the same. The model we propose implements a mapping between the lexical item have and the continuous semantic space schematized in Figures 1 and 2. Instead of articulatory movement goals, peaks of neural activation correspond to semantic interpretations. One contribution of the paper is to offer a neurocognitive process explanation of the results described above: contextual modulation of the timecourse and outcome of lexical interpretation, and individual variation in the magnitude of this modulation. More generally, we offer an explanation for the intuition that meaning is experienced as discrete (e.g., adjacency or possession), despite evidence for a continuous substrate. Such apparent discreteness, we propose, arises from metastability in a continuous space defined by neurocognitive variables. In addition to capturing existing results, simulations from the model also generate a novel empirical prediction, which we test with an experiment combining self-paced reading and acceptability judgments. We intend our model to be general, using the English lexical item have as a test case or proof of concept. We propose that, while individual lexical items vary in the relevant semantic dimensions, as well as in the details of their coupling to those dimensions, the basic architecture and mechanisms of our model do not vary.
The remainder of the paper is structured as follows. Section 2 describes the structure of the model and the results of the simulations. Section 3 describes the design of the experiment, and Section 4 reports the experimental results. Section 5 discusses theoretical implications of the study and new predictions that remain to be tested.
2 Dynamic neural model of lexical meaning
2.1 Model structure
The model consists of two dynamic neural fields (DNFs) and one dynamic node. One DNF governs interpretation on the semantic dimension of control asymmetry (CA), and the other DNF governs interpretation on the dimension of connectedness (conn). The node corresponds to the lexical item have. Activation of the node follows simple linear dynamics, given in Eq. 1.
The rate of change of activation is negatively related to current activation , defining a dynamical system with a point attractor at . represents external input to the node, e.g. from perception or intention, and represents normally distributed noise weighted by a parameter . Thus, when there is external input to the node, the node’s activation is attracted to . When there is no input, activation is attracted to 0. is a time constant, with higher values corresponding to slower rates of evolution. In the simulations reported below, is set to , and is set to . We set the magnitude of external input depending on the condition being simulated.
Activation in each of the two DNFs unfolds according to Eq. 2.
Activation is defined for each neuron representing the relevant semantic dimension at each moment in time . The range of is set to 100, such that each neuron in the field is maximally sensitive (“tuned”) to a particular percentage of the maximum conceivable value of that dimension. For instance, when the neuron in the conn DNF is active, this corresponds to an interpretation of “ of maximum conceivable connectedness”. Activation in the DNF has a point attractor at . The resting level is assumed to be below zero for all neurons, by convention at . We set the time constant for the DNFs to . Each field input , , and is represented as a separate Gaussian distribution of the form
where controls the amplitude of the input, controls the position of the input in the field, and controls the width of the input distribution. In the simulations reported below, we set the amplitude of external input depending on the condition being simulated. represents input from the have node, whose amplitude is defined straightforwardly as a linear function of the activation of the node: . This is a simplification relative to most DFT models, where the amplitude of input from a node to a field would be a more complex (sigmoidal) function of node activation. Our motivation for eliminating the sigmoidal gating function on node-to-field input comes from work on lexical neighborhood effects on articulation, where non-selected lexical items (nodes) exert some influence on phonetic planning (DNFs) (Stern and Shaw, 2023b). This issue is largely orthogonal to the present study because there is only one node in the model. We chose to use the same node-field coupling dynamics as in Stern and Shaw (2023b) for the sake of simplicity and consistency. The same results could likely be achieved with a relatively shallow or “soft” sigmoidal gating function.
The have node is coupled to the conn DNF with a wide distribution ( = 40) positioned at the center of the field ( = 50), as seen in Figure 3. When the have node becomes active, it sends input to the entire conn field, consistent with an analysis of polysemy for the lexical semantics of have. However, the center of the conn field is favored, consistent with the fact that, all else equal, have tends to evoke an interpretation of alienable possession (intermediate connectedness).
represents input from each DNF to the other. Via this mechanism, patterns of activation in one DNF evoke corresponding patterns of activation in the other DNF. This mechanism implements the “inverted U” pattern described in Section 1 (Figure 2). In particular, as seen in Figure 4, activation consistent with high control asymmetry evokes activation consistent with intermediate connectedness (and vice versa; note the double-sided arrows), and activation consistent with low control asymmetry evokes activation consistent with both low connectedness and high connectedness (and vice versa).
The amplitude of each input from one DNF to the other is given by
The term on the right, , defines the basic magnitude of . The magnitude of input from each neuron in the field which is sending input (of size ) to the field receiving input is determined by its activation above resting level , weighted by a parameter , set to 0.3 in our simulations. This magnitude is further weighted by , a nonlinear (quartic) function of the distance from to the center of the sending distribution , divided by the width of the sending distribution . Neurons within one of contribute substantially to , while neurons exceeding one from contribute exponentially less. This value is summed for all neurons within the sending field, and then weighted by the term on the left, . When the maximum activation within the sending distribution (ranging from to ) is much greater than the maximum activation in the receiving distribution, this term approaches 1, so . However, as the maximum activation in the receiving distribution approaches (or exceeds) the maximum activation in the sending distribution, the weighting term approaches 0, reducing or eliminating . In this way, the sending field cannot increase activation in the receiving field beyond its own maximum activation, preventing an infinite positive feedback loop. The centers and widths of each distribution are given in Table 1.
| distribution | ||
| low CA | ||
| high CA | ||
| low conn | ||
| mid conn | ||
| high conn |
Within-field lateral interaction between neurons is defined by an interaction kernel :
Each neuron which exceeds an activation threshold contributes activation to other neurons as a function of their distance within the field . As seen in Figure 5, interaction is excitatory (weighted by , set here to ) for nearby neurons (defined by , set here to ) and inhibitory (weighted by , set here to ) for more distant neurons (defined by , set here to ). , set here to 2, contributes global inhibition from each above-threshold neuron.
Lateral excitation helps to stabilize activation peaks (which correspond to semantic interpretations), and lateral inhibition prevents runaway expansion of activation peaks. Crucially, we set the parameters of the interaction kernel such that only a single peak can form at a time in a given field for the range of input amplitudes under consideration in our simulations, defining selection dynamics.
As seen in Eq. 6 and Figure 6, the activation threshold for interaction is given by a sigmoidal function , where (set here to ) controls the steepness of the threshold:
By convention, the threshold is set to so that lateral interaction kicks in only when activation approaches . Finally, noise in field activation is simulated by adding normally distributed random values weighted by , set here to .
2.2 Model simulations
In this section, we use the model to simulate the interpretation of sentences conveying adjacency meanings of have in two contexts: (1) following an adjacency interpretation of have, and (2) following a possession interpretation of have. Each simulation consists of three phases, summarized in Table 2. In phase 1, the model receives external inputs (all with amplitude ) which drive interpretation of the context sentence. In the adjacency context, these inputs correspond to low CA and low conn; the have node also receives input. In the possession context, external inputs also excite the have node and the low side of the CA field, but the conn field receives input corresponding to high connectedness, i.e., inalienable possession. Phase 1 is the only phase which differs between conditions. In phase 2, all inputs are removed. This corresponds to the time between interpreting the context sentence and interpreting the target sentence. In phase 3, the model receives external inputs corresponding to the adjacency have target sentence. These inputs are identical in both conditions: have and low CA. Phases 1 and 3 each run for 90 timesteps, which we found to be enough time for an activation peak to stabilize in both DNFs. Phase 2 runs for 20 timesteps, which we found to be enough time for both activation peaks to fall below the interaction threshold (), but not enough time for the fields to return fully to the resting level. Each simulation runs for a total of 200 timesteps.
| Phase 1: context | Phase 2: no input | Phase 3: target | |
| adjacency context | low CA low conn have | - | low CA have |
| possession context | low CA high conn have | - | low CA have |
Examples of simulated interpretation in each context are displayed in Figure 7. In phase 1 (up to timestep 90) the have node becomes active, and begins to send input to the conn field. The CA field quickly forms a stabilized peak corresponding to an interpretation of low control asymmetry in both conditions. Due to field coupling, this CA peak sends input to the conn field on both the low and high ends. In the adjacency context, an activation peak corresponding to low conn stabilizes and suppresses the rest of the field via lateral inhibition. In the possession context, the winning activation peak is on the high side of the field. During phase 2 (from timesteps 91–110), all activation decreases towards resting level. Activation of the have node returns almost back to its resting level. In the fields, activation falls below the interaction threshold (), but does not return fully to the resting level. At timestep 111 (the beginning of phase 3), field activation is still highly non-uniform, reflecting residual activation from phase 1. In the CA field, another peak on the low side of the field forms in both conditions. In the conn field, the location of the peak differs by condition, despite the fact that the phase 3 inputs are identical between conditions. In the adjacency context, phase 3 (target) processing stabilizes on another interpretation of low conn; in the possession context, the stabilized peak corresponds to an interpretation of high conn. In this way, the evolution of activation is shaped by the preceding state of the system.
The examples in Figure 7 were selected in order to demonstrate the effect of preceding context (phase 1 evolution) on target sentence interpretation (phase 3 evolution). However, not every simulated run of the model exhibits the same effect. The presence of noise in the model introduces a stochastic influence on the location of field stabilization in each simulation. In order to examine the robustness of the contextual modulation effect, we simulated 1000 instances of interpretation in each of the two conditions. As seen in Figure 8, there is a bimodal distribution of interpretations in both contexts. That is, for each context, an interpretation of either adjacency or possession for the target sentence was possible. However, the likelihood of each interpretation was influenced by context. In the adjacency context, low conn (adjacency) interpretations were more likely; in the possession context, high conn (possession) interpretations were more likely. Thus, while context does not completely determine the course of field evolution, it exerts a strong enough influence to be observable over many simulations.
How do we relate these simulated results to the observed acceptability results described in Section 1? Field stabilization at the low end of the conn field corresponds to an adjacency interpretation, which is consistent with the two inanimate arguments of have in the target sentence. Thus, a low conn peak should correlate with a higher acceptability rating. On the other hand, a high conn peak, corresponding to a possession interpretation, is inconsistent with the two nouns in the target sentence, which cannot easily be construed to be in a possession relationship. Thus, a high conn peak should correlate with a lower acceptability rating. In order to transform the field stabilization results into simulated acceptability results, we therefore reverse the value of the location of field stabilization using the following equation
where ranges from to . As seen in Figure 9, simulated acceptability was higher in the adjacency context relative to the possession context, consistent with existing results.
Existing empirical results have also demonstrated that individual variation in the magnitude of the by-context difference in acceptability is predicted by AQ scores, such that individuals with higher AQ scores show a reduced influence of context. We model individual variation in AQ by varying the parameter , which controls the magnitude of field coupling. Stronger field coupling is consistent with a greater degree of system-level expectations. In other words, given some interpretation on one semantic dimension (e.g., control asymmetry (CA)), individuals can vary in the degree to which they expect a corresponding interpretation on a related semantic dimension (e.g., connectedness (conn)). We posit that individuals with higher AQ scores are more influenced by system-level expectations, i.e. stronger field coupling. With stronger system-level expectations, the processing system is more rigid, and less influenced by real-time signals. This is consistent with existing findings relating AQ scores to linguistic behavior. For instance, individuals with higher AQ scores show greater compensation for coarticulation in speech perception (Yu, 2010). In addition, higher AQ individuals are less sensitive to phonetic duration when assigning judgments of prosodic prominence (Bishop, 2016). Both sets of results suggest that higher AQ individuals rely more on system-level expectations (e.g., expected cooccurrence of phonetic signals), and less on the real-time signal itself (e.g., the phonetic duration of a perceived word).
In order to apply this hypothesis to the case of contextual influence on the interpretation of have, we varied from 0.3 to 0.7 in steps of 0.1, and at each level, we ran 1000 simulations in each condition. As seen in Figure 10, higher values of corresponded with a reduction in the difference in acceptability between conditions. In other words, stronger field coupling reduced the magnitude of the contextual modulation effect. This is because input from the low CA distribution equally favors the low conn and high conn distributions. When this input is stronger, the lingering asymmetry between these distributions from phase 1 is reduced more quickly, reducing the influence of context on interpretation in phase 3.
By modeling the timecourse of individual instances of semantic interpretation, we can examine the relationship between interpretation content (the location of field stabilization) and interpretation time (the timestep of field stabilization) for individual simulations. As seen in Figure 11, this relationship was modulated by context. In the adjacency context, a negative correlation was observed: faster response times corresponded with higher acceptability ratings. This is because the adjacency context primed a more acceptable interpretation: when the field stabilized at a lower conn value (higher acceptability), this stabilization was facilitated by the context. In the possession context, a positive correlation was observed: faster response times corresponded with lower acceptability ratings. This is because the interpretation primed by the possession context was ultimately a less acceptable one.
To date, studies of contextual facilitation of have interpretation have not simultaneously collected data regarding acceptability and processing time. Below we report an experiment combining acceptability judgments and self-paced reading. The purpose of the experiment is twofold: (a) to replicate previous results regarding contextual facilitation of adjacency readings of have, as well as individual variation in the magnitude of this effect indexed by the AQ, and (b) to test the new model prediction represented in Figure 11.
3 Experiment design
3.1 Participants
56 adults participated in the experiment (ages 20-30; 32 women, 21 men, 3 nonbinary). All participants self-reported that they were native monolingual speakers of American English, and that they had no history of speech, language, hearing, or reading impairment. Participants were recruited through Prolific (www.prolific.com). Before beginning the experiment, participants provided informed consent under Yale University IRB #2000033871.
3.2 Materials
Each experimental stimulus consisted of a pair of sentences (a context sentence followed by a target sentence) conjoined by and. Every target sentence was designed to convey an adjacency interpretation of have, as in 3.2.
. …the oak tree has a skateboard that is red.
It is difficult to construe a possession interpretation of 3.2 because oak trees do not typically possess skateboards (whether alienably or inalienably). Every target sentence had the form “the [noun1] has a [noun2] that is [adj]”. [noun2] in the target sentence is the critical word at which an interpretation of have can be construed, since after reading [noun2], the participant has read have and both of its arguments. The relative clause “that is [adj]” was included as a spillover region. There were ten target sentences, each of which was preceded by two different context sentences, for a total of 20 experimental stimuli. Each context sentence conveyed either an adjacency reading or an inalienable possession reading of have, as in Table 3.
| adjacency | possession |
| The maple tree has a plastic box behind it and… | The maple tree has a branch that is thick and… |
Every context sentence had the structure “the [noun1] has a [noun2] [modifier]”. [modifier] was either a prepositional phrase (in the adjacency condition) or a “that is [adj]” phrase (in the possession condition). [noun1] was identical between the two conditions in each set, and always contrasted saliently with [noun1] in the target sentence in order to increase overall felicitousness. In both conditions, [noun2] in the context was semantically unrelated to [noun2] in the target sentence. Moreover, the first phoneme in [noun2] in the context was always different from the first phoneme of [noun2] in the target sentence, in order to minimize confounds from phonological priming. All nouns were inanimate in order to maximize the availability of adjacency interpretations. 60 filler stimuli were also included. 20 of the fillers were of the same form as the experimental stimuli, but with contexts that used verbs other than have to convey an adjacency reading (10 stimuli) or a possession reading (10 stimuli). 40 of the fillers were completely unrelated to the experimental stimuli: 20 conveyed interpretations of circumstantial metonymy (e.g., “the grilled cheese at Table 6 ordered another coffee:), and 20 conveyed non-metonymous counterparts (e.g., “the customer at Table 6 ordered another coffee”). This yielded a total of 80 stimuli (20 experimental + 60 fillers).
In order to examine effects of individual variation in communicative context sensitivity, participants completed the Autism-Spectrum Quotient (AQ: Baron-Cohen et al., 2001). The AQ consists of 50 statements (e.g., “I prefer to do things the same way over and over again”). The participant responds to each statement by selecting one of four options: “definitely disagree”, “slightly disagree”, “slightly agree”, “definitely agree”.
3.3 Procedure
Participants were instructed to complete the experiment in a quiet room where they would be free from distractions for up to an hour. To begin each trial, participants clicked on a button at the top of the screen with the words “Click here to begin the next trial”. Then, the first word of the stimulus appeared in the center of the screen. Participants pressed the spacebar to advance to the next word. Participants were instructed to read as quickly as possible while making sure to comprehend what they were reading. After participants advanced past the last word, they were prompted by the following instruction to give an acceptability rating of the entire stimulus: “How likely would you be to say this sentence, or hear this sentence from another native speaker of English?”. Participants gave their response on a seven-point Likert scale (labeled “very unnatural” on the left end and “very natural” on the right) by clicking on the corresponding button at the bottom of the screen.
The experiment began with four practice trials unrelated to the experimental stimuli. Then, each stimulus was presented twice, for a total of 160 trials per subject. Each of the two blocks was pseudo-randomized such that no two consecutive trials were from the same condition or the same stimulus set. The presentation order of the two blocks was counter-balanced between participants. After completing the combined self-paced reading and acceptability judgment task, participants completed the AQ. The entire procedure was conducted in the same session in Gorilla (Anwyl-Irvine et al., 2020). The session lasted approximately 30-40 minutes.
3.4 Data processing
For the analysis of acceptability ratings, trials with acceptability response times greater than 2.5 standard deviations from the participant’s mean were removed. This led to the exclusion of 194 trials (2.20%). Before plotting the rating data, raw ratings were z-scored by participant in order to abstract away from idiosyncratic rating styles (e.g., staying towards the ends or towards the center of the scale). In statistical models, this was accomplished via random effects by participant. For the analysis of word reading times, reading times less than 120 ms or greater than 2000 ms were removed (210 trials, 2.39%). For one participant, 38% of their trials were removed according to these criteria. This participant’s data was subsequently completely excluded from analysis, leaving a total of 2098 experimental trials from 55 participants for analysis.
Reading times were log-transformed in order to approximate a normal distribution. In order to assess the effects of control variables on reading time, a linear mixed effects model was fit to the log-transformed RT (logRT) of all words with fixed effects of word length (in characters), trial number, and the preceding word’s logRT (all scaled and centered), a random intercept by participant, and random slopes for all predictors by participant. The model results are displayed in Table 4. All three control variables were found to significantly affect logRT: words took longer to read when they had more characters or when the preceding word took longer to read. Words were read faster when the word came later in the experiment. Thus, rather than analyzing raw logRT, we analyze the residuals of logRT after being regressed, by participant, on the three control predictors described above. In other words, we analyze the variance not predicted by the three control predictors.
| Estimate | Std. Error | df | t value | p value | |
| char_length | |||||
| trial_num | |||||
| prec_logRT |
Each AQ response was initially coded on a four-step Likert scale where “definitely disagree” = 1 and “definitely agree” = 4. For half of the questions on the AQ, agreement signifies greater autistic traits and disagreement signifies reduced autistic traits; this is reversed for the other half of questions. For questions where disagreement signified greater autistic traits, we reversed the numerical response so that higher values always corresponded to greater autistic traits. Thus, individual AQ scores could range from 50 (very low autistic traits) to 200 (very high autistic traits).
4 Experiment results
4.1 Acceptability
In order to assess the effect of context on acceptability, we fit nested linear mixed effects models to acceptability ratings using the lme4 package (Bates et al., 2015) in R (R Core Team, 2021). All models included random intercepts by item and by participant, and random slopes for context by item and by participant. Model comparison revealed that a control fixed factor for trial number (scaled and centered) significantly improved model fit over a baseline model that only included random effects (, ). Additionally including the experimental fixed factor of context (treatment coded; reference level = possession) significantly improved model fit over the control model (, ). The results of the full model are displayed in Table 5. Trial number significantly decreased acceptability ratings, such that ratings generally decreased over the course of the experiment. Regarding the experimental factor of interest, ratings were significantly higher in the adjacency context relative to the possession context.
| Estimate | Std. Error | df | t value | p value | |
| trial number | |||||
| context=adjacency |
Figure 12 displays mean z-scored (by-participant) acceptability ratings in each context. Both means are below 0, suggesting that the experimental stimuli were generally less acceptable than the filler stimuli, consistent with the general markedness of adjacency interpretations of have. However, consistent with the results of the regression model, ratings were higher in the adjacency context relative to the possession context.
Recall from Section 2 that in the model, context affected interpretation not by shifting the overall distribution of acceptability, but by increasing the likelihood of a distinct, higher-acceptability interpretation. We compare this simulated behavior to the measured behavior by examining the distribution of ratings in each context. Figure 13 demonstrates that, consistent with the simulations, the increase in mean acceptability in the adjacency context relative to the possession context was not accomplished via on overall shift in the distribution of ratings. For instance, the number of intermediate ratings 3 and 4 was largely unaffected by context. Rather, the increase in mean acceptability occurred because of a decrease in low ratings 1 and 2, and an increase in high ratings 5 and 6.
In order to assess whether individual variation in the magnitude of contextual facilitation is predicted by AQ scores, we plot the by-participant random slopes for the effect of context against AQ score in Figure 14. A Spearman test confirms a negative correlation (, ): subjects with higher AQ scores showed a smaller effect of context, i.e. reduced contextual facilitation of adjacency have sentences.
4.2 Relationship between acceptability and reading time
Next, we turn to the by-trial relationship between acceptability ratings and reading times. Recall that the model simulations predicted this relationship to be modified by context: a negative correlation in the adjacency context, and a positive correlation in the possession context. To test this prediction, we fit nested linear mixed effects models to the summed residualized reading times of the critical word [noun2] and each of the words in the spillover region “that is [adj]”. All models included random intercepts by item and by participant. Random slopes were not included because they led to model convergence issues and were also not of theoretical interest in this case.
Adding a fixed factor of context did not lead to a significant improvement in fit over a baseline model with only random effects (, ). However, adding a fixed factor for acceptability rating (z-scored by subject) did lead to a significant improvement in fit over the baseline model (, ). Adding a fixed factor for context back in did not lead to an improvement in fit over the model with only a fixed factor for acceptability (, ). Importantly, however, adding an interaction term did lead to a significant improvement in fit over the model with only fixed factors for main effects (, ). We report the results of the full model with both main effects and their interaction in Table 6. We observed a significant negative interaction between context and acceptability, such that the relationship between acceptability and reading time was more negative in the adjacency context relative to the possession context. This result, also observable in Figure 15, is consistent with the model prediction described in Section 2.2.
| Estimate | Std. Error | df | t value | p value | |
| context=adjacency | |||||
| acceptability | |||||
| context:acceptability |
4.3 Discussion
The acceptability results confirm previously reported results regarding contextual facilitation of adjacency have interpretations. The acceptability of adjacency have sentences was improved following a context sentence that also conveyed an adjacency have interpretation, relative to a context sentence that conveyed a possession have interpretation. Moreover, individual variation in the magnitude of this improvement was predicted by AQ scores: individuals with higher AQ scores (lower context-sensitivity) showed reduced contextual facilitation. Finally, consistent with the novel prediction of the model simulations, the by-trial relationship between reading time and acceptability evidenced context modulation. The relationship was more negative in the adjacency context relative to the possession context.
5 General discussion & conclusion
5.1 Summary & discussion
We have argued for a dynamic neural model of lexical meaning and demonstrated its behavior using the English lexical item have as a test case. The cognitive basis of the model is a continuous meaning space with two parameters: control asymmetry and connectedness. Interpretations of “possession” and “adjacency” associated with have result from the neural dynamics governing interpretation along those dimensions. The apparent discreteness of the different interpretations is epiphenomenal, resulting from the property of metastability in the neural dynamics.333The concept of metastability has previously been invoked in cognitive science most frequently to explain phenomena in visual perception, observed from both behaviorial (e.g., Chaudhuri and Glaser, 1991) and brain measures (e.g., Sterzer et al., 2009). Metastability has also been invoked to explain semantic ambiguity in language comprehension (Wildgen, 1995). Our model represents a mathematically explicit demonstration of the role of metastability in language comprehension.
In the model, lexical meaning is a coupling pattern between a neural node representing the lexical item and dynamic neural fields (DNFs) governing interpretation on continuous semantic dimensions. Interpretation occurs in time as activation of the lexical node causes peaks of activation in the semantic DNFs. The locations of the peaks in feature space correspond to the content of the interpretation. Dependencies between semantic dimensions, schematized in Figures 1 and 2, are modeled as coupling patterns between DNFs.
Simulations from the model captured known empirical effects. In particular, the interpretation evoked by English have was influenced by a preceding interpretation. An adjacency interpretation was more likely following another adjacency interpretation, and a possession interpretation was more likely following another possession interpretation. This occurred in the model because activation states persist in time, continuing to bias interpretation until they return to a resting state.444The model parameter , which controls the rate of field evolution, modulates the temporal extent of contextual bias on interpretation. Model simulations also exhibited covariation between the magnitude of contextual modulation and the strength of coupling between DNFs. Stronger coupling makes DNFs more resistant to the effects of context. In other words, there is a stronger influence of long-term knowledge, relative to immediate context, on lexical interpretation. Motivated by other observations regarding individual variation in speech behavior (described in Section 2.2), we related DNF coupling strength to the Autism-Spectrum Quotient (AQ).
Finally, model simulations generated a novel prediction: an interaction between context, acceptability, and response time. Specifically, the prediction was that in an adjacency context, more acceptable (adjacency) interpretations would be reached more quickly, while in a possession context, more acceptable interpretations would be reached more slowly. This prediction was empirically confirmed by the results of a novel experiment combining self-paced reading and acceptability judgments. Results showed an interaction between the effects of context and acceptability on response time on a trial by trial basis, such that the relationship between acceptability and response time was more negative in the adjacency context relative to the possession context. That model simulations were able to generate this empirical prediction demonstrates an important advantage of the DFT framework. In DFT, both time and cognitive features are modeled as continuous dimensions. This allows the generation of precise quantitative predictions regarding the relationship between the timecourse and outcome of cognitive processing.555In the domain of phonetics, for example, the spatial target of an articulatory movement can be related to the time it takes to generate that target (Roon and Gafos, 2016; Stern and Shaw, 2023b). As demonstrated in this study, when DNFs represent semantic dimensions, we can relate the content of a semantic interpretation to the time it takes to arrive at that interpretation. This allows us to move beyond the assumption that less acceptable sentences should always take longer to process, and thus that acceptability and response time should always be negatively correlated.
5.2 Novel predictions of our approach
Our proposal that individual variation in the strength of coupling between DNFs is indexed by the AQ generates testable predictions, including in domains unrelated to linguistic meaning. For example, in an experimental task that requires learning associations between object color (one DNF) and object shape (another DNF), individuals with higher AQ scores are predicted to show greater surprisal effects (e.g., slowed response times) when encountering an object that violates the learned associations. Modeling learned dependencies between dimensions as DNF coupling may also shed light on other issues in linguistics. For example, phonological inventories (possible sounds in a language) and phonotactic constraints (possible sound sequences in a language) may be explained as the result of coupling between DNFs representing phonetic dimensions. Under this view, DNF coupling patterns are an important component of language-specific knowledge, and thus an important dimension of variation between speakers of different languages. Among speakers of the same language, processes conditioned by phonological knowledge, such as accent and perceptual illusions (Davidson and Shaw, 2012; Dupoux et al., 1999; Hallé and Best, 2007; Kabak and Idsardi, 2007), are predicted to covary with AQ scores.
5.3 Other perspectives
Before concluding, we briefly discuss the relationship of our proposal to other perspectives on lexical meaning. The proposal that lexical items map to continuous semantic spaces contrasts with models of lexical meaning based on discrete concept representations (e.g., Dell, 1986; Levelt et al., 1999; Roelofs, 1997; Stella et al., 2024). In these models, lexical meaning is formalized as a set of weighted connections between an activation node representing the lexical (e.g., CAT) and activation nodes representing discrete concept representations (e.g., ALIVE, FURRY, etc.). These models capture the intuition of the discreteness of semantic interpretations (e.g., adjacency or possession), but do not explain the evidence for a continuous substrate described in Section 1.3. In Section 2, we explained how our model fills that gap by showing how it reconciles intuitions of discreteness with evidence for continuity.
Our proposal also contrasts with distributed representations of lexical meaning, e.g., distributed parallel processing networks (e.g., Kawamoto, 1993; Rumelhart et al., 1986), or vectors describing cooccurrence patterns in written language corpora (e.g., Baayen et al., 2019; Landauer and Dumais, 1997; Mikolov et al., 2013). These models capture important facts related to the semantic similarity of lexical items, e.g., priming effects on reaction times (Kawamoto, 1993). But in contrast to our current effort, their focus is not on understanding the dimensions themselves on which lexical meanings vary. As a result, their models consist of a very large number of dimensions, each of which is not interpretable on its own. For example, in a vector-based model, the dimensionality of a word’s semantic representation is equal to the number of words in the training corpus. In a parallel distributed processing model, the dimensionality equals the number of connections between processing units. Unlike the model we proposed here, each dimension of description, on its own, is not intended to be meaningful with respect to cognition.
By contrast, in our model, interpretations are modeled as peaks of activation in DNFs representing cognitively viable semantic dimensions. This model thus allows us to understand the relatedness of the various interpretations of a polysemous lexical item as proximity in the semantically parameterized space represented by DNFs. It further allows us to understand the difference between polysemy and homophony as a difference in the proximity of possible activation peaks in the space defined by the semantic DNFs to which the lexical node is coupled. Possible activation peaks are more proximal in the case of polysemy and more distal in the case of homophony. This difference arises from differences in the coupling patterns between lexical nodes and semantic DNFs. This kind of solution demonstrates an advantage of a low-dimensional “localist” framework like DFT.666This perspective also has the potential to shed light on observed differences in the timecourse of processing polysemous and homophonous words. Different interpretations of homophonous words tend to inhibit each other, leading to slowed response times, while different interpretations of polysemous words tend to prime each other, leading to faster response times (Frisson, 2015; Klepousniotou and Baum, 2007; Klepousniotou et al., 2012; MacGregor et al., 2015; Rodd et al., 2002). These facts may be explained as a result of the basic dynamics of lateral interaction within DNFs. As described in Section 2, neurons which are proximal in a DNF tend to excite each other, while those which are more distal tend to inhibit each other. These are the dynamics which allow the formation of activation peaks. Neural recordings consistent with these dynamics have been observed in rhesus monkey motor cortex (Georgopoulos et al., 1986) and cat visual cortex (Jancke et al., 1999). The difference between polyseme and homophone processing can be seen as an effect of metric feature distance on target-distractor interaction, analogous to effects observed in the domains of eye saccades (Kopecz and Schöner, 1995), manual reaching movements (Erlhagen and Schöner, 2002), and speech articulation (Tilsen, 2009).
5.4 Conclusion
Using a single theoretical framework (DFT) and mathematics of description (differential equations) allows explicit integration across the cognitive and sensory-motor domains. Previous DFT modeling work has linked neural representations of conceptual structure (Jackendoff, 2002) with visual perception (Grieben and Schöner, 2022) and visual search behavior (Sabinasz et al., 2023). Explicit coupling between abstract cognitive processes and sensation/movement has been termed “grounding” of cognition (e.g., Sabinasz et al., 2023; Sabinasz and Schöner, 2023). As described in Section 1.4, existing DFT models of speech and language have focused on the sensory-motor domain, with DNFs representing phonetic dimensions of articulatory movement and auditory perception. In this paper we have extended DFT to linguistic meaning through the lens of lexical polysemy, a cognitive domain that appears relatively distinct from sensation and behavior. Our use of the same theoretical framework and mathematics as these previous models paves the way for explicitly coupling the cognitive and sensory-motor aspects of language. This would represent a significant step towards a complete neurocognitive model of language from meaning to form: a grounded model of linguistic cognition.
6 Data availability
Data (experimental and simulated) and scripts for analysis and simulation are available on OSF at https://osf.io/3bue9/
7 Funding sources
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.